Professor Very Busy needs to allocate time next week to include time for office hours. He needs to forecast the number of students who will seek appointments. He has gathered the following data:

Week #Students
6 Weeks ago 83
5 Weeks ago 110
4 Weeks ago 95
3 Weeks ago 80
2 Weeks ago 65
Last Week 50

  
What is this week's forecast using exponential smoothing with alpha = .2, if the forecast for two weeks ago was 90?
Answer #1

Using the simple exponential smoothing method

F(Last week) = 0.2*Actual(Two weeks ago) + (1-0.2)*F(Two weeks ago)

= 0.2*65 + 0.8*90

= 13 + 72

= 85

F(This week) = 0.2*50 + 0.8*85

= 78